OpenStudy (anonymous):

The Eighth term is 75 and the 20th term is 39 A.)Find the first term and the common difference

OpenStudy (raden):

use the formula : a_n = a1 + (n-1)d a8 = 75 -----> a1 + 7d = 75 a20 = 39 ----> a1 + 19d = 39 from both equation above, you can eliminate d, to get a1

OpenStudy (anonymous):

how exactly is "d" eliminated?

OpenStudy (whpalmer4):

you've got two equations: \[a_1 + 7d = 75\]\[a_1 + 19d = 39\]-------------- subtract down the columns and you'll get an equation involving only \(d\). Solve it. Then plug the value you found back into either of the equations to solve for \(a_1\). Done.

OpenStudy (whpalmer4):

You could also eliminate \(d\) first, but that would require an extra step of multiplying one or both equations by a factor so that the \(d\) "column" has equal coefficients like the \(a_1\) column already has...

OpenStudy (anonymous):

so like a sytem of equations?

OpenStudy (whpalmer4):

Yes, that's exactly what you're doing — solving a system of two equations in two unknowns.

OpenStudy (anonymous):

sweet, thanks alot

OpenStudy (whpalmer4):

you're welcome!